# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v4

@@ D.1. Fixed Train and Test Missingness Fractions
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 30. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 31-48. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across ML- and DL-based imputers and all our datasets.
@@ D.2. Variable Train and Test Missingness Rates
Other fairness metrics under variable train set missingness. Figures 52, 53, and 54 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
Other fairness metrics under variable test set missingness. Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.

### v5

@@ D.1. Fixed Train and Test Missingness Fractions
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 33. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 34-51. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across ML- and DL-based imputers and all our datasets.
@@ D.2. Variable Train and Test Missingness Rates
Other fairness metrics under variable train set missingness. Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
Other fairness metrics under variable test set missingness. Figures 58, 59, and 60 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.